OpenStudy (anonymous):

in what base b is the equation 53*15 = 732 valid? All three of the numbers in the equation are base b, and b E z.

OpenStudy (anonymous):

B^53 B^15 = B^732 ?? B^68 = B^732 doesnt make sense...

OpenStudy (anonymous):

B=1 if anthing :)

OpenStudy (anonymous):

if b is the base, then \[53 = 5 \times b ^{1}+3 \times b ^{0}\]

OpenStudy (anonymous):

\[15 = 1 \times b ^{1} + 5 b ^{0}\]

OpenStudy (anonymous):

\[732 = 7 \times b ^{2}+3 \times b ^{1} + 2 \times b ^{0}\]

OpenStudy (anonymous):

equate the two. ie, (5b+3)(b+5) = 7b^2+3b +2 solve for b

OpenStudy (anonymous):

Im Hotep55's answer is incorrect., b cannot be 1.

OpenStudy (anonymous):

aii thx

OpenStudy (anonymous):

np